OpenStudy (anonymous):

does (2x+15)(x-2)=2x^2+13x+15

OpenStudy (anonymous):

15(2) = 30 ..NOT 15...so I would have to say no

OpenStudy (anonymous):

I just got that...

OpenStudy (anonymous):

you typoed or else its wrong lol

OpenStudy (anonymous):

how do you get thirteen?

OpenStudy (anonymous):

No I just got that it was wrong lol

OpenStudy (anonymous):

id be 15x-4x if anything and even that's bad....

OpenStudy (anonymous):

FOIL the thing out dude :)

OpenStudy (anonymous):

I don't think that's right.

OpenStudy (anonymous):

I tried. I hate algebra lol

OpenStudy (anonymous):

(2x+15)(x-2) 2x(x) + 2x(-2) +15(x) +15(-2)

OpenStudy (anonymous):

2x^2 -4x +15x -30 2x^2 +11x -30

OpenStudy (anonymous):

then wouldn't it be using 10 and 3?

OpenStudy (anonymous):

would what be using 10 and 3 ? what would make it true you mean?

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

what are the factor of 15? 1,15 3,5 thats your options right there if you want integer answers.

OpenStudy (anonymous):

that's wrong too

OpenStudy (anonymous):

well what I said

OpenStudy (anonymous):

15+1 = 16 3+5 =8...so you dont get a right answer lol

OpenStudy (anonymous):

Well what my teacher taught us to multiply the end numbers and factor the number from that

OpenStudy (anonymous):

i hate shortcuts as answers, unless you understand the mechanics of what is going on in the problem; shortcuts just confuse the issue.

OpenStudy (anonymous):

distribute the equations thru...

OpenStudy (anonymous):

tell me: what does it mean to distribute?

OpenStudy (anonymous):

a(b+c) = ?? a(b) + a(c) right?

OpenStudy (anonymous):

I know what it means. This review for us. I just don't remember any of it.

OpenStudy (anonymous):

did i just "distribute" that a thru the paranthesis?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(f+l) (m+n) distribute the first into the last there: (f+l)(m) + (f+l)(n) Now distribute again :) f(m) + l(m) + f(n) + l(n)

OpenStudy (anonymous):

that is HOW you solve the problem....the shortcut is FOIL or whatever acronym they feel like using...

OpenStudy (anonymous):

when you forget the shortcuts, use the actual process ;)

OpenStudy (anonymous):

ok, thanks lol